### v2

# Better Aligned with Survey Respondents or Training Data? Unveiling Political Leanings of LLMs on U.S. Supreme Court Cases

Abstract
The increased adoption of Large Language Models (LLMs) and their potential to shape public opinion have sparked interest in assessing these models’ political leanings. Building on previous research that compared LLMs and human opinions and observed political bias in system responses, we take a step further to investigate the underlying causes of such biases by empirically examining how the values and biases embedded in training corpora shape model outputs. Specifically, we propose a method to quantitatively evaluate political leanings embedded in the large pretraining corpora. Subsequently we investigate to whom are the LLMs’ political leanings more aligned with, their pretrainig corpora or the surveyed human opinions. As a case study, we focus on probing the political leanings of LLMs in 32 U.S. Supreme Court cases, addressing contentious topics such as abortion and voting rights. Our findings reveal that LLMs strongly reflect the political leanings in their training data, and no strong correlation is observed with their alignment to human opinions as expressed in surveys. These results underscore the importance of responsible curation of training data and the need for robust evaluation metrics to ensure LLMs’ alignment with human-centered values.
[…]
Better Aligned with Survey Respondents or Training Data?
Unveiling Political Leanings of LLMs on U.S. Supreme Court Cases
Shanshan Xu1, Santosh T.Y.S.S1,Yanai Elazar2,3 Quirin Vogel1, Barbara Plank4,5, Matthias Grabmair1 1Technical University of Munich, Germany 2Allen Institute for AI 3University of Washington 4IT University of Copenhagen, Denmark 5LMU Munich & Munich Center for Machine Learning (MCML)
1 Introduction
Values and opinions expressed by large language models (LLMs) such as ChatGPT OpenAI (2023) have the potential to influence public perceptions and shape societal views. As such, it is essential to evaluate the political leanings of LLMs. Previous research (e.g., Perez et al. 2023; Hartmann et al. 2023) found a misalignment between LLM outputs and human opinions, and several identified a left-leaning bias of LLMs.
Figure 1: Assessing the political leanings of LLMs, and comparing it with that in their training data, and of human respondents.
LLMs derive their knowledge primarily from their pre-training data, which are typically composed of internet text. However, the structure of online content regularly skews toward overrepresenting specific viewpoints Galeazzi et al. (2024), often introducing biases into these models. Such biases can then propagate through model outputs, reinforcing existing societal biases Feng et al. (2023). Despite increasing awareness of these issues, to the best of our knowledge, no prior work has shown how to quantitatively analyze political bias in the content of large pre-training corpora and subsequently assess how well LLM outputs align with human opinions versus the biases in training data. To address this gap, we propose a pipeline to retrieve relevant documents from the pretraining corpora, then evaluate the political leanings expressed in these documents, and subsequently assess the alignment of political leanings in pretraining corpora with the responses generated by the LLM.
As a case study, we focus on US Supreme Court cases, which frequently address contentious and politically charged issues, such as death penalty, abortion, same-sex marriage, and voting rights, making them strong indicators of political leanings. Leveraging the SCOPE Jessee et al. (2022) [^1] Jessee et al. 2022 has not named the dataset. Hereafter, we refer to the dataset as SCOPE: Supreme COurt Case Political Evaluation. survey data on US Supreme Court cases from political studies, this paper examines the political leanings of eight LLMs and five open-source pre-training corpora, comparing them to human survey responses and Supreme Court rulings [^2] We will release the code and data upon publication.. The main contributions of our work are threefold:
  • We conduct a quantitative analysis of political bias in large pre-training corpora by examining the political stance of the documents in the corpora.
  • We compare LLMs’ alignment with both surveyed human opinions and with their pre-training corpora (as illustrated in Fig 1).
  • Our empirical findings indicate that LLMs exhibit significant alignment with their training corpora, yet we do not find strong alignment with human opinions. This highlights the critical need for responsible data curation and rigorous evaluation of LLM alignment.
2 Background
Evaluating LLM-Human Alignment
Recent research has increasingly focused on probing LLMs political opinions. Most approaches typically follow a two-stage process: (1) assessing an LLM’s political stance on specific topics, and (2) measuring how closely its responses align with human opinions. A common strategy for evaluating LLM opinions involves using political orientation tests (e.g., Political Compass Test333www.politicalcompass.org/test, as in Röttger et al. 2024; Feng et al. 2023) or survey questionnaires (e.g., PewResearch ATP444www.pewresearch.org/writing-survey-questions/, as in Santurkar et al. 2023). To quantify the alignment between human and LLM responses, prior work typically measures the similarity of their opinion distributions using either (a) distance-based metrics—such as the 1-Wasserstein distance Santurkar et al. (2023); Sanders et al. (2023) and Jensen–Shannon divergence Durmus et al. (2024)—or (b) statistical analyses, including Cohen’s Kappa Argyle et al. (2023); Hwang et al. (2023) and Pearson correlation coefficients Movva et al. (2024). We refer to Ma et al. 2024 for an extensive survey of methods in this area. In this work, we use SCOPE to probe LLMs’ political opinions because it offers several advantages over the above-mentioned political surveys used in previous studies: The cases in SCOPE are selected by experts, ensuring that they address the most important and publicly salient legal topics. Experts carefully word the questions and response options to be understandable to the general public. Moreover, political experts have annotated each case with thoughtfully chosen keywords, which facilitate our retrieval of relevant documents from large pretraining corpora, as detailed in Sec 4.1.3.
LLMs and their pretraining corpra
Existing studies have explored the impact of biases in training corpora on LLM behavior, primarily through second-stage controlled training setups such as continual pretraining Feng et al. (2023); Chalkidis and Brandl (2024). While continual pretraining can offer valuable insights into the causal links between training data and model outputs, these studies rarely applied to study LLMs’ behavior based on initial pretraining phase, where biases are fundamentally embedded. Additionally, it is also computationally expensive to conduct such extensive pre-training experiments on initial phase. An alternative strategy involves investigating the correlation between biases in training corpora and those in model outputs Seshadri et al. (2024). Previously this approach has been under exploited, primarily due to the limited accessibility of large-scale pretraining datasets. Many commercial LLM providers (e.g., GPT-4 by OpenAI 2023 and Claude by Anthropic 2023) disclose minimal information about their training sets, not even corpus size or data source. With the growing call in the academic community for transparency and accessibility of LLM pretraining data Pistilli et al. (2023); McMillan-Major et al. (2024), several organizations have begun to make large-scale pretraining datasets publicly available, including RedPajama (Weber et al., 2024) and Dolma Soldaini et al. (2024). These initiatives are complemented by the development of APIs and analytical tooling platforms, such as WIMDB Elazar et al. (2024), which facilitate comprehensive analysis of the corpora. In this paper, we leverage WIMDB to analyze the political leanings in five publicly accessible corpora and subsequently evaluate how these leanings correlate with the outputs generated by various LLMs.
3 Experimental Setup
3.1 Dataset
In political science, researchers often estimate individuals’ or groups’ political preferences and ideological positions by analyzing observable behaviors, such as voting patterns and survey responses Martin and Quinn (2002); Ho et al. (2008). For example, Jessee et al. (2022) created SCOPE to gather respondents’ views on Supreme Court decisions. By comparing collected survey responses with the Court’s voting record, they demonstrated that the Court has adopted a more conservative stance than the general U.S. public. In this study, we use SCOPE Jessee et al. (2022) to prompt various LLMs to assess their political leanings and subsequently compare their alignment with surveyed human opinions and political leanings in their training data.
The SCOPE dataset comprises of 32 cases, each represented by a binary-choice question asking respondents to express their views on the Court’s ruling as either supportive (pro) or opposing (opp). Fig 2 provides an example of a survey question. Tab 2 in App C lists all 32 cases along with their corresponding legal topics in the SCOPE dataset. For each case, between 1,500 and 2,158 respondents indicate whether they are pro or opp regarding the Court’s decision. Additionally, SCOPE captures each respondent’s self-identified political ideology, enabling the categorization of participants into self-identified Democrats or Republicans. Tab 2 in App C showcases the distribution of choices {pro, opp} among the overall surveyed respondents, as well as within the self-identified Democratic and Republican respondents. Further descriptive statistics on respondents’ backgrounds are available in the original study Jessee et al. (2022).
Figure 2: An example case from the SCOPE Jessee et al. (2022) dataset. In this case, 53.6% of the surveyed respondents agreed with the court’s decision (option 1). When broken down by party affiliation, 77.4% of self-identified Republicans and 29% of self-identified Democrats supported the court’s decision.
Company Model Short Name Model Full ID Size Pretraining Data
OpenAI GPT-4o GPT-4o Unknown Unknown
Allen AI OLMo-sft OLMo-7B-SFT-hf 7B Dolma
OLMo-instruct OLMo-7B-0724-Instruct-hf 7B Dolma
Google Gemma gemma-7b-it 7B Unknown
Meta Llama3-8b Llama-3-8B-Instruct 8B RedPajama*
Llama3-70b Llama-3-70B-Instruct 70B RedPajama*
Big Science T0 T0 11B C4*
BLOOMZ BLOOMZ-7b1 7B OSCAR*, The Pile*
Table 1: Details of evaluated LLMs, along with the dataset that the model trained on. * signifies that the model was not trained exactly on this dataset, due to filtering, using additional data, or the original data being private.
3.2 Evaluated LLMs
We evaluate eight models that have been fine-tuned for instruction following and conversational abilities. This includes seven open-source models: Gemma-7b-it Team et al. (2024), Llama-3-8B-Instruct, Llama-70B-Instruct Dubey et al. (2024), OLMo-7B-Instruct, OLMo-7B-SFT Groeneveld et al. (2024), BLOOMZ Muennighoff et al. (2022), and T0 Sanh et al. (2021), as well as one closed-source model, GPT-4o OpenAI (2023). Details about these models can be found in Tab 1. Further implementation details are discussed in App A.
3.3 Pretraining Corpora
Tab 1 lists the corresponding pretraining corpora (when available) of the LLMs we investigated in this work. It is important to note that among the various LLM-corpora pairs we consider, only the OLMo-SFT and OLMo-Instruct models were trained directly on the pretraining corpus Dolma Soldaini et al. (2024). While for all other pairs, the LLMs may not have been trained exactly on the versions of the corpora we consider, due to factors such as filtering, or inclusion of additional data Elazar et al. (2024). Despite these discrepancies, we treat the documented corpora as reasonable proxies for analysis, as they represent the closest publicly available approximations of the actual training data for these models [^5] All models examined in this paper have undergone post-training, such as Supervised or Instruction Fine-Tuning, which may also influence the opinions in models outputs. However, prior research Feng et al. (2023) suggests that the shift introduced by post-training is relatively small. We also explored the correlation between LLMs’ political leanings and that in their post-training data, but did not observe any significant correlation. Further discussions can be found in App G..
4 Methodology
We employ a three-stage process to examine LLMs’ political leanings and further compare their alignment with surveyed human opinions and their pretraining corpora, when available. First, we introduce how we assess the political leanings of different entities [^6] We use entity to refer to either a group of surveyed respondents, Supreme Court justices, LLM-generated responses, or content within the training data.. Next, we measure the political leanings alignment among them in Sec 4.2. Finally, we conduct significance tests to determine whether the observed differences in LLM alignment with different entities are statistically significant in Sec 4.3.
Preference Distributions
In our study, we assess the political leanings of various entities by analyzing their preference distributions on SCOPE. We define preference distributions on a survey as follows: consider a survey consisting of a series of questions denoted as $\mathcal{Q}=\left\{q_{i}\right\}_{i=1}^{m}$ , where each question $q_{i}$ offers $n$ possible choices $\left\{a_{j}\right\}_{j=1}^{n}$. For our binary questionnaire $n=2$, and the generalization to more choices are straight-forward. For an entity $k$, we define its political preference distribution $D_{k}\in\mathbb{R}^{m\times n}$ as:
$D_{k}^{ij}=p_{k}(a_{i}|q_{j})\in[0,1],$
[…]
4.1 Extracting the Preference Distributions
In this section, we outline the methodology used to extract the preference distribution of various entities. We divide the entities to three categories - the humans $D_{H}$, the LLMs $D_{M}$, and the pretraining corpora $D_{C}$.
4.1.1 The Humans
Under the category of the humans, we consider the preference distribution of four entities: $D_{H}=\{D_{\mathrm{pub}},D_{\mathrm{dem}},D_{\mathrm{rep}},D_{\mathrm{court}}\}$. Here, $D_{\mathrm{pub}}$ represents the preference distribution of the overall surveyed respondents, while $D_{\mathrm{dem}}$ and $D_{\mathrm{rep}}$ correspond to surveyed self-identified Democrat and Republican respondents, respectively; $D_{\mathrm{court}}$ represents the preference distribution of the Court. All preference distributions are Bernoulli, with the respective parameter estimated from the data. For the $D_{\mathrm{court}}$, we fetch the judges’ votes from the Supreme Court Database Spaeth et al. (2024) [^7] http://scdb.wustl.edu/ , and then calculate $D_{\mathrm{court}}$ as the ratio of justices who agree (pro) / dissent (opo) with the majority decision. For $D_{pub},D_{\mathrm{dem}}$ and $D_{\mathrm{rep}}$ we calculate them as the ratios of {pro} versus {opo} to the court’s decisions among the respondents based on data retrieved from SCOPE.
Figure 3: Extracting the Preference Distributions of the Pretraining Corpora.
4.1.2 The LLMs
Under the category of the LLMs $D_{M}$, we probe the political preferences of eight LLMs as listed in Tab 1. Following Scherrer et al. 2023, for each survey case in SCOPE, we created six different prompt templates, and we then sample five responses from each of the six prompt variations from the LLMs at a temperature setting of 1, yielding in a total of 30 responses per case per model. Example prompt templates and detailed prompt creation process can be found in Fig 6 in App B.
To map the LLM-generated answers to one of the given choice options, we employed an iterative, rule-based matching pipeline, as explained in App B. The preference distribution, denoted as $D_{m}=p_{m}(a_{j}\mid q_{i})$, reflects the ratio of support versus opposition to the court’s decision across the 30 generated responses for model $m$ on case $q_{i}$, illustrated in Fig 6 in App B.
4.1.3 The Training Data
[…]
(i) Relevant Document Retrieval
For each of the 32 cases $q_{i}$ in the SCOPE survey, we compile a set of keywords $K_{i}$ to retrieve relevant documents $T_{i}$ from the pretraining corpora using the WIMDB API Elazar et al. (2024)- a tool designed to facilitate analysis of large-scale pretraining corpora. For example, in the case Baze v. Rees [^8] Baze v. Rees, 553 U.S. 35 (2008), addresses whether lethal injection for executions was constitutional or not., we use keywords such as [lethal injection; capital punishment; Baze; Rees], retrieving 206 documents from the Dolma corpus. Further details on keyword selection and retrieval statistics for each case are provided in App C. Additionally, an example of a retrieved document is included in App I.
(ii) Stance Score Evaluation
We use Llama3-70B to assess the political stance $s_{i}^{j}$ of each retrieved document $t_{i}^{j}\in T_{i}$. The model is prompted to evaluate the document’s level of support for the court’s decision on a Likert scale from 1 (strongly against) to 5 (strongly supportive). If a document is unrelated to the case’s political issue, the model is instructed to return ‘Not related.’
(iii) Preference Distribution Estimation
To quantify the political leaning of each case $q_{i}$, we first compute the average stance score $S_{i}=\frac{1}{m}\sum_{j=1}^{m}s_{i}^{j}$ where $s_{i}^{j}$ denotes the stance score of a retrieved document assigned by Llama3-70B on a Likert scale ranging from 1 (strong opposition) to 5 (strong support). To facilitate probabilistic interpretation, we transform $S_{i}$ from its original Likert scale to a probability measure $P_{i}$, which represents the likelihood that the document supports the court’s decision.
Quality Assessment of Stance Detection To evaluate the reliability of Llama370B’s stance detection, we manually annotated a randomly selected sample of 80 retrieved documents. We measure the agreement between human and model labels using Spearman’s rank correlationSpearman (1904). The overall Spearman’s $\rho$ is 0.68, indicating good alignment between Llama370B and human annotators. App C offers details on the quality assessment process. To evaluate the robustness of our document retrieval method, we performed a bootstrapping analysis by iteratively excluding 20% of retrieved documents. This procedure revealed no significant shifts in measured political leanings (see App C for methodological details). Although differences in keyword selection may affect document retrieval and thereby influence corpus-level political stance estimates, our findings demonstrate that results are resilient to changes in the retrieved documents.
4.2 Measuring the LLMs Alignments
We use Pearson correlation to measure the alignment over distribution pairs of different respondents/entities. We define alignment between two preference distribution $D_{1}$ and $D_{2}$ on a set of questions $\mathcal{Q}$ as:
$\rho\left(D_{1},D_{2}\right)=\operatorname{CoRR}\left(D_{1},D_{2}\right)\,,$
where $\operatorname{CoRR}$ calculates the Pearson correlation coefficients when averaged across questions. The $p$-value associated to the Pearson coefficient quantifies statistical significance Kowalski (1972).
4.3 Testing for Significance of Stronger Alignments
Given an LLM $D_{m}$ and two human groups $D_{\mathrm{dem}}$ and $D_{\mathrm{rep}}$, we compute the alignments $\rho(D_{m},D_{\mathrm{dem}})$ and $\rho(D_{m},D_{\mathrm{rep}})$. To determine whether $D_{m}$ aligns more strongly with $D_{\mathrm{dem}}$ than $D_{\mathrm{rep}}$, a direct comparison of $\rho(D_{m},D_{\mathrm{dem}})$ and $\rho(D_{m},D_{\mathrm{rep}})$ is insufficient. This is because both correlations are derived from the same dataset, meaning they are statistically dependent. Consequently, standard significance tests for independent correlations fail to account for the covariance between $\rho(D_{m},D_{\mathrm{dem}})$ and $\rho(D_{m},D_{\mathrm{rep}})$, potentially overestimating or underestimating the significance of their difference. To address this, we use a variation of Williams test (Williams, 1959), which evaluates the significance of differences in dependent correlations Steiger (1980). This test has been widely adopted for comparing the performance of machine translation and text summarization metrics (Mathur et al., 2020; Deutsch et al., 2021; Graham and Baldwin, 2014). In essence, it tests whether the population correlation between $D_{1}$ and $D_{3}$ equals the population correlation between $D_{2}$ and $D_{3}$, where the test-statistic is given by:
$t_{n-3}=\frac{\left(\rho_{12}-\rho_{13}\right)\sqrt{(n-1)\left(1+\rho_{12}\right)}}{\sqrt{2K\frac{(n-1)}{(n-3)}+\frac{\left(\rho_{12}+\rho_{13}\right)^{2}}{4}\left(1-\rho_{23}\right)^{3}}},$
where $\rho_{ij}$ is the correlation between $D_{i}$ and $D_{j},n$ (i.e., $\rho_{ij}=\operatorname{CoRR}(D_{i},D_{j})$) is the size of the population, and $K$ can be computed as:
$K=1-\rho_{12}^{2}-\rho_{13}^{2}-\rho_{23}^{2}+2\rho_{12}\rho_{13}\rho_{23}\,.$
Figure 4: Pearson Alignment. Cell $(i,j)$ represents the Pearson correlation $\rho$ of LLM $i$ to entity $j$. * shows p-value < 0.05, **∗ ∗ shows p-value < 0.001.
Figure 5: The result of Williams significance tests, in each subfigure, where a colored cell in row $i$ (named on y-axis), col $j$ (named on x-axis) indicates that the LLM $m$ correlates significantly higher with entity $i$ than entity $j$, at a significance level of $0.05$.
5 Results and Anyalsis
This section presents the results and analysis of our experiments. Our investigation on the alignment of LLMs can be formed into two key questions: (1) Is there a statistically significant correlation between the preference distribution of LLM $m$ and the entity $E_{1}$? (2) Given $m$, $E_{1}$, and $E_{2}$, is the correlation between $(m,E_{1})$ significantly stronger than that between $(m,E_{2})$? To address the first question, we applied Pearson correlation to quantify the alignment between LLMs and different entities. Fig 4 presents a heatmap depicting the Pearson correlation coefficients ($\rho$-values) between LLMs, surveyed human opinions ($D_{H}$), and pretraining corpora ($D_{C}$). For the second question, we employed the Williams test to assess whether the observed differences between correlation pairs are statistically significant, as shown in Fig 5. Due to space constraints, our discussion highlights the Williams test results for six selected LLMs. A full overview of all LLMs’ results is provided in Fig 12 in App H. We make the following observations:
Takeaway 1: LLMs are primarily aligned with their pretraining data, but not with surveyed human opinions.
Fig 4 illustrates the alignment of various LLMs with surveyed human opinions alongside their pretraining corpora, when applicable. Notably, both versions of OLMo-instruct ($\rho=0.63$) and OLMo-sft ($\rho=0.57$) demonstrate a significant correlation with Dolma (highly significant $p<0.001$), which is precisely the pretraining corpus utilized for their training. Similarly, although the correlation is not as statistically significant, the T0 model exhibits the strongest correlation with its pretraining corpus, C4, compared to the other five training corpora.
In contrast to the observed trends in monolingual LLMs, the multilingual BLOOMZ exhibits no statistically significant correlation with the aforementioned three pretraining corpora. We hypothesize that its political preference patterns may stem from exposure to non-English languages in training data, which includes different distribution of political views from the English-only corpora we evaluated. This aligns with prior research showing that multilingual models trained on diverse language data can develop unpredictable moral and political biases Hämmerl et al. (2023).
Furthermore, all LLMs, with the exception of Bloomz and T0, display a significant positive correlation with the three training corpora: Dolma, RedPajama, and Oscar. This alignment may stem from the similar political leanings in these corpora and the models trained on them [^9] Fig 11 in App F presents the alignments between different training corpora and surveyed human opinions. . In contrast, our findings indicate that there are generally no significant alignments between the LLMs’ outputs and surveyed human opinions. The only LLMs that do not follow this trend are LLama3-70b and T0, which we will discuss further in Takeaway 3.
Takeaway 2: Significance testing confirms LLM’s alignment to their pretraining data is stronger than to humans.
Fig 5 illustrates the results of the Williams tests conducted on various pairs of alignments. As demonstrated in subfigures Fig 5 a), b), and c), GPT4-o, OLMo-sft, and OLMo-instruct consistently exhibit a significantly stronger alignment with the training corpora (Dolma, RedPajama, Oscar) than with human groups, $p<0.05$. This finding corresponds to the orange cluster in Fig 4, confirming that these LLMs have a stronger alignment to the pretraining data than to the surveyed human opinions.
Takeaway 3: Correlation numbers alone are not enough.
To address the question, “With which entity $E_{k}$ is model $M$ most aligned?”, it is crucial to not only compare the strength (correlation coefficient $\rho$) and significance (p-value) of each correlation $(m,E)$; but also to determine whether the correlation between $(m,E_{1})$ (statistically) significantly differs from that between $(m,E_{2})$. As discussed in Sec 4.3, the dependencies of these distributions imply that a higher correlation coefficient, $\rho(m,E_{1})>\rho(m,E_{2})$, does not necessarily indicate that model $M$ is more aligned with $E_{1}$ than with $E_{2}$, even for small $p$-values. Therefore, a significance test is needed to ascertain whether model $M$ is significantly more aligned with $E_{1}$ compared to $E_{2}$, or if the observed differences in $\rho$ values are attributable to random variation. For example, as illustrated in Fig 4, the preference distribution of LLama3-70B exhibits significant correlations ($p<0.05$) with both its pretraining corpus, RedPajama ($\rho=0.53$), and the $E_{dem}$ (surveyed democratic respondents, ($\rho=0.48$). However, according to the Williams test results in Fig 5(d), the correlation between LLaMA3-70B and RedPajama is not significantly different from its correlation with $E_{dem}$ the Democratic respondents, indicating that the observed difference in the correlation pearson coefficient $\rho$ could be due to statistical noise.
Similarly, the Pearson correlation results in Fig 4 indicate that T0 exhibits a significant correlation only with $E_{\text{pub}}$ (surveyed human opinions), while no significant correlations are observed with other entities. At first glance, this might suggest that T0 is most aligned with human opinions among all entities. However, the significance test results in Fig 5(e) reveal inconsistencies. While correlation between $(T0,E_{\text{pub}})$ is significantly stronger than the correlation between $(T0,E_{\text{court}})$, no such significant differences are found with other entities, such as with $E_{\text{dem}}$ or any of the training corpora. This means that we can only conclude that T0 aligns more closely with surveyed human opinions $E_{\text{pub}}$ than with the court $E_{\text{court}}$ , but we cannot determine whether its alignment with $E_{\text{pub}}$ is significantly stronger than its alignment with other entities, even though there are great differences in $\rho$-values observed in Fig 4.
These two examples from LLaMA3-70B and T0 underscore the limitations of evaluating LLM alignment based solely on correlation values and highlight the importance of significance testing.
6 Implications and Future Directions
Training Data Curation
Our empirical results indicate that LLMs closely reflect the political leanings present in their training data, raising concerns given the lack of transparency and accountability in the data curation process. Historians Harari (2024) compare this process to the canonization of religious text, in which a group of religious authorities decides which works to include or exclude, subsequently shaping the evolution of beliefs and societal norms. Similarly, a small group of AI engineers determine which sources are deemed “trustworthy” and which are classified as “harmful”, ultimately shaping the epistemic landscape of AI-generated knowledge. To mitigate these issues, the AI community can adopt “datasheets”Gebru et al. (2021), which is widely used in the community benchmark datasets. The datasheets should document key metadata, including data sources, filtering methodologies, and known biases or limitations. Policymakers, in turn, should establish legal frameworks mandating independent audits and risk assessments of training data curation.
Public Discourse Framework
Our research reveals that most LLMs exhibit alignment with their training corpora, yet not necessarily with the surveyed human opinions. Nonetheless, in the public discourse framework, attributing human characteristics to AI - known as anthropomorphizing — seems to be quite natural. This tendency may lead to an over-reliance on AI, as users might confuse AI-generated responses for human beings, leading to excessive trust PAIR (2019). Further, anthropomorphism can obscure accountability, shifting the responsibility away from developers and onto the LLM itself. Recent studies McCoy et al. (2024) suggest moving away from anthropomorphic and advocate for a reframing of public dialogue in alternative conceptual frameworks, such as viewing LLMs as simulation systems of the integration of diverse perspectives in their training data Shanahan et al. (2023). In conclusion, fostering a clear public understanding of the distinctions between AI and human beings is essential for a more responsible engagement with AI technologies.
7 Conclusion
We introduced a pipeline to investigate political leaning in the pretraining corpora, which allows us to compare the LLMs’ political leaning not only with surveyed human opinions but also the political leanings embedded in their pretraining corpora. By examining LLMs’ stances on political issues derived from U.S. Supreme Court cases, our results reveal a significant alignment between the models and their training corpora, yet no similarly strong alignment with human opinions as survyed is found. We call on the AI community to explore methods for rigorous evaluation of LLM alignment, and advocate for more transparent and collaborative strategies in curating training data for LLM.
Limitations
Multi-choice Format
Our work probes LLMs’ political views using questions from a public opinion survey, requiring LLMs to answer in a binary-choice format. However, the methodology laid out in this article does not rely on the binary format. Correlation coefficients, the Williams test and the Jensen–Shannon divergence immediately generalize to more refined analysis of political biases, such as continuous distributions, multiple-choice formats or clusterings. Recent research Röttger et al. (2024) indeed suggests that such constrained formats may not accurately reflect real-world LLM usage, where users tend to talk in open-ended discussions on controversial topics Ouyang et al. (2023). They also found in unconstrained settings, LLMs may respond differently than when restricted to a fixed set of options. We leave this question to future analysis. Furthermore, we point out that in certain real-world applications, such as voting assistants Chalkidis (2024), often necessitate LLMs to function within a binary or multiple-choice framework.
Partisan Aggregation in Political Alignment Analysis
Our analysis compares LLMs’ political leanings to human survey responses aggregated by partisan groups, such as Democrats and Republicans. However, this approach has inherent limitations. Political opinions on controversial issues can resist strict partisan categorization, as individuals within the same party do not always align neatly with partisan divisions, as individuals within the same party may hold diverse or even opposing views. Recent research has highlighted the pluralism of human opinions and proposed incorporating fine-grained human values into AI systems Plank (2022); Xu et al. (2024); Sorensen et al. (2024). Future research could explore LLMs’ response uncertainty—using metrics such as entropy or confidence scores across multiple generations—to assess whether these models capture the ambiguity of opinions on contentious topics. We call for more work to contribute to aligning LLMs with pluralistic human values.
U.S.-Centric Perspectives
While the expert-chosen cases within SCOPE address contentious issues and serve as strong indicators of political orientation, the framework is not without its limitations. Notably, akin to other political surveys employed in recent LLM evaluation studies (e.g. ANES in Bisbee et al. 2024), SCOPE is based on U.S. centric public opinion data and focuses on the American partisan political ideology. This emphasis constrains its applicability when assessing LLMs that have been trained on multilingual or globally diverse datasets, as showed in our experiment results on the BLOOMZ model. Despite these limitations, we propose a method that enables comparisons between the alignment of LLMs with the surveyed human opinions and their pretraining corpora, thus enabling flexibility across various ideological frameworks or questionnaires. We encourage future research to adopt our approach on alternative ideological theories and political surveys. This will contribute to a more comprehensive understanding of LLMs’ political positioning.
Acknowledgments
We thank Rashid Haddad, Lukas Gosch, Valentin Hoffman, Max Prior and Yang Li for the helpful discussions.
Information about use of AI assistants
In the preparation of this work, the authors utilized ChatGPT-4o and Grammarly to correct grammatical errors and improve the coherence of the manuscript. Before the submission, the authors conducted a thorough review and made necessary edits to the content, taking full responsibility for the final version of the text.
Appendix A Implementation Details
We downloaded the OLMo-SFT, OLMo-Instruct, LLama3-7b models, BLOOMZ and T0 from HuggingFace Hub Wolf et al. (2020) and ran the downloaded LLMs on an A100 GPU. We accessed the other models through the DeepInfra API. We use default generation parameters from the ‘transformers’ library, except for temperature. We set temperature to 1 to when probing LLMs views on SCOPE cases. When using LLama3-70 to detect the stance score of training documents, we set temperature to 0 to reduce variation to a minimum . We collected all GPT responses in November 2024.
Figure 6: LLM prompt
Appendix B LLM Response Collection
Fig 6 demonstrates how we prompt the LLMs for responses. Prior research has shown that LLMs are sensitive to the prompt format and the sequence of answer options Webson and Pavlick (2022), and they may display inconsistencies in their responses Elazar et al. (2021). To mitigate these issues, we implemented three variations of prompts, following Scherrer et al. (2023). We also randomize the order of the answer choices within each format, producing six unique prompt forms. 7(c) demonstrates the prompts we used to query the LLMs’ political preference.
Figure 7: Prompts Used to Query Political Preference
(a) Question Template: AB
(b) Question Template: Repeat
(c) Question Template: Compare
Mapping LLM Response to Preferences
To map LLM generated sequences of tokens to actions (i.e., opinion preference), we use an iterative, rule-based matching pipeline in the following order:
1. Check for exact matches (i.e., check for exact overlaps with the desired answer, such as "A" or "Yes")
2. Check for normalized matches (e.g. "A)" or "YES"). For the few unmatched sequences, we manually coded the actions.
Court Decision % agreeing with decision
docket case_name Keywords Full Sample Public Reps Dems
19-123 Fulton v. City of Philadelphia PA Gay Adoption; religious; Fulton; Philadelphia Conservative 52.2% 65.6% 38.9%
19-1257A
Brnovich v. Democratic National
Committee I
Provisional Ballot; precinct; Brnovich;
Democratic National Committee
Conservative 49.1% 67.7% 33.9%
19-1257B
Brnovich v. Democratic National
Committee II
Ballot Harvesting; third party; Brnovich;
Democratic National Committee
Conservative 50.0% 75.2% 29.2%
19-251
Americans for Prosperity Foundation
v. Becerra
Donors; information; Americans for
Prosperity Foundation; Becerra
Conservative 40.0% 56.1% 25.5%
20-255 Mahanoy Area School District v. B.L.
School; Free Speech; punish; off campus;
Mahanoy Area School; B.L.
Conservative 70.5% 80.2% 63.3%
18-1259 Jones v. Mississippi Juvenile; criminal; life sentence; Mississippi; Jones Conservative 29.4% 36.0% 22.0%
19-783 Van Buren v. United States
Government; databases; authorized; access;
Van Buren; United States
Conservative 31.9% 31.8% 31.2%
20-512 National Collegiate Athletic Association v. Alston college athletes; compensation; limit; NCAA; Alston Liberal 49.9% 39.7% 58.1%
20A87 Roman Catholic Diocese of Brooklyn v. Cuomo COVID; religious gathering; church; Cuomo Conservative 53.6% 77.4% 29.0%
20-107 Cedar Point Nursery v. Hassid Unions; enter; private property; Cedar Point Nursery; Hassid Conservative 51.6% 70.6% 34.4%
19-422 Collins v. Mnuchin Federal Agencies; Collins; Mnuchin Conservative 45.5% 50.1% 39.9%
20-18 Lange v. California police; warrant; private property; Lange; California Liberal 52.4% 41.2% 60.0%
17-1618 Bostock v. Clayton County, Georgia Fire; Gay Employees; Bostock; Clayton Liberal 83.3% 74.6% 90.4%
18-107
R.G. & G.R. Harris Funeral Homes Inc. v.
Equal Employment Opportunity Commission
Fire; Transgender Employees;
Equal Employment Opportunity Commission
Liberal 78.8% 68.6% 87.2%
18-587
Department of Homeland Security v.
Regents of the University of California
Deferred Action for Childhood Arrivals;
Department of Homeland Security; University of California
Liberal 61.0% 30.4% 85.5%
18-1195 Espinoza v. Montana Department of Revenue Scholarship; taxpayer; religious school; Espinoza; Montana Conservative 63.1% 76.6% 52.3%
19-431
Little Sisters of the Poor Saints Peter
and Paul Home v. Pennsylvania
Contraceptives; health insurance;
Little Sisters; Pennsylvania
Conservative 52.7% 70.4% 33.3%
18-1323 June Medical Services, LLC v. Russo
Abortion; admitting privileges; constitutional
rights; June Medical Services; Russo
Liberal 56.9% 37.3% 73.6%
19-635
Trump v. Deutsche Bank AG and
Trump v. Mazars USA, LLP
Trump; taxes; Mazars; President; Congress;
Deutsche Bank
Liberal 60.9% 30.9% 84.5%
19-715 Trump v. Vance Trump; taxes; President; state prosecutors; Vance Liberal 61.3% 28.0% 85.5%
19-7 Seila Law, LLC v. CFPB CFPB; independent; power; Seila Law Conservative 43.6% 69.4% 20.8%
19-465 Chiafalo v. Washington Electoral College; Chiafalo; Washington Liberal 61.4% 59.5% 65.0%
08-1224 U.S. v. Comstock Sex Offenders; prison; Comstock Conservative 54.5% 52.6% 50.1%
08-1521 McDonald v. Chicago Gun Control; government; ban; possession; McDonald; Chicago Conservative 71.4% 92.8% 56.6%
08-472 Salazar v. Buono Religious Symbols; public land; Salazar; Buono Conservative 62.1% 85.4% 43.0%
07-1428 Ricci v. DeStefano Affirmative Action; racial diversity; Ricci; DeStefano Conservative 89.6% 95.7% 81.8%
07-21 Crawford v. Marion County Voter; photo identification; Crawford; Marion County Conservative 81.6% 92.7% 75.3%
08-205 Citizens United v. FEC
prohibit; corporations; political campaign;
Citizens United; Federal Election Commission
Conservative 44.7% 67.9% 27.4%
07-5439 Baze v. Rees Lethal injection; capital punishment; Baze; Rees Conservative 78.8% 93.2% 70.9%
05-908 Parents Involved v. Seattle Race; Schools; admission; diversity; Parents Involved; Seattle Conservative 84.8% 98.1% 74.9%
05-380 Gonzales v. Carhart
Partial Birth Abortion; federal government;
ban; Gonzales; Carhart
Conservative 55.0% 78.6% 37.8%
05-184 Hamdan v. Rumsfeld Guantanamo Bay; trial; suspected terrorist; Hamdan; Rumsfeld Liberal 30.4% 9.9% 53.2%
Table 2: The distribution of choices among the respondents, together with the Keywords used to retrieve relevant documents from the pretraining data
Appendix C Keyword List
We define the keywords for each case as [keyword 1, keyword 2, plaintiff, defendant], with the two keywords derived from Jesse’s original dataset. We manually adjusted some keywords as necessary to refine the search scope. Including the names of the parties enhances the precision of document retrieval, because in the U.S., cases are typically cited using the names of the parties involved in the format “plaintiff v. defendant". When acronyms or abbreviations are commonly used, we manually edit the party names for better retrieval result; for example, we use NCAA instead of the full name “National Collegiate Athletic Association". The complete list of keywords of all cases are available in Tab 2. An example of a retrieved document is provided in App I.
Appendix D Relevant Documents Retrieval
We used the WIMBD API Elazar et al. (2024) to retrieve documents based on defined keywords. Due to the API and token limitations of LLama3, we retrieved only documents with word counts below this threshold. Fig 10 displays the distribution of document lengths, showing that most contain fewer than 4,000 words. Tab 3 provides additional statistics such as the number of documents matching the keywords in the Dolma dataset (documents matched) and the subset we fetched (those with fewer than 4,000 words, documents fetched)
Figure 8: Prompt used to evaluate the stance scores of the retrieved documents.
Appendix E Quality Assessment of Stance Detection
To evaluate the quality of LLaMA3-70B’s stance detection, we conducted a two-round quality assessment. In the first round, we randomly sampled 20 documents from the retrieved relevant documents. Two annotators independently labeled the documents: Annotator 1, a research assistant who is a native English speaker and a U.S. citizen, and Annotator 2, the first author of this paper. The annotation process followed the exact template used to prompt LLaMA3-70B, as shown in Fig 8. The inter-annotator agreement, measured by Spearman’s $\rho$, was 0.76. The Spearman’s $\rho$ between Annotator 1 and LLaMA3-70B’s labels was 0.7. In the second round, Annotator 1 labeled an additional 40 documents. The overall Spearman’s $\rho$ between Annotator 1 and LLaMA3-70B’s labels across all 60 documents was 0.68. Based on this, we consider the alignment between LLaMA3-70B’s outputs and human annotations to be strong.
Figure 9: Bootstrapped sample means and their 95% confidence intervals for each docket. Each bar represents the average stance score for a given case docket, while the error bars denote the 5th and 95th percentiles of the bootstrap distribution (based on repeatedly sampling 80% of the data).
Bootstrap Resampling
We applied a bootstrap resampling procedure to assess the robustness of political stance score estimation. For each of the 32 cases in SCOPE, we generated 100 bootstrap samples by randomly subsampling 80% of its retrieved documents’ stance scores. The mean score was computed for each subsample, creating a bootstrap distribution of means. We derived 95% confidence intervals (CIs) using the percentile method, with bounds defined by this distribution’s 5th and 95th percentiles. The sample mean (calculated on the full dataset) and its CI bounds were recorded for all dockets. As shown in Fig 9 , all sample means lie within their respective CIs, confirming the reliability of our estimates and quantifying their variability.
docket Survey wave # doc fetched avg. length doc fetched avg. stance score # doc matched avg. length doc matched
19-123 2021 59 1,346 3.40 113 11,111
19-1257A 2021 124 1,369 2.27 165 5,414
19-1257B 2021 384 1,148 4.42 441 4,900
19-251 2021 338 909 4.13 396 2,784
20-255 2021 520 978 4.23 577 1,739
18-1259 2021 150 2,256 4.01 2,139 49,361
19-783 2021 257 1,573 4.27 1,205 24,803
20-512 2021 602 1,292 1.46 683 3,830
20A87 2021 96 1,848 3.57 206 12,722
20-107 2021 282 993 1.47 300 1,592
19-422 2021 277 1,862 3.19 825 22,763
20-18 2021 57 1,323 4.76 488 122,176
17-1618 2020 87 1,513 4.75 102 4,025
18-107 2020 405 1,246 4.87 562 5,803
18-587 2020 2,005 1,047 4.31 2,216 2,311
18-1195 2020 224 1,318 4.02 292 4,835
19-431 2020 699 1,204 3.99 908 6,311
18-1323 2020 185 1,712 3.60 253 7,368
19-635 2020 878 1,181 1.56 1,214 6,384
19-715 2020 456 1,216 1.51 693 12,316
19-7 2020 1,047 1,159 4.19 1,205 2,920
19-465 2020 814 1,040 3.63 895 2,015
08-1224 2010 316 1,389 2.21 565 11,954
08-1521 2010 1,570 2,380 4.47 5,903 32,346
08-472 2010 108 1,165 2.62 172 13,837
07-1428 2010 72 1,884 3.12 141 10,650
07-21 2010 727 1,366 3.08 945 5,139
08-205 2010 187 2,613 2.89 731 20,631
07-5439 2010 927 1,458 2.25 1,360 5,799
05-908 2010 471 1,802 2.73 974 15,752
05-380 2010 402 2,012 3.53 983 12,542
05-184 2010 56 2,419 3.17 176 20,370
Table 3: Descriptive statistics of the documents retrieved from the Dolma dataset.
Figure 10: Distribution of length of all the matched documents.
Figure 11: Pearson Alignment. Cell $(i,j)$ represents the Pearson correlation $\rho$ of LLM $i$ to entity $j$. * shows *p* value < 0.05, **∗ ∗ shows p-value < 0.001.
Appendix F Corpora-Human Alignment
Fig 11 presents the alignments between different training corpora and surveyed human opinions. The political leanings of these pretraining corpora appear to be quite similar; however, they differ from those of the human respondents surveyed. Further, among the 5 corpora, DOLMa, RedPajama and OSCAR high correlation to each other. They are less correlated to C4 and the Pile, which might be due to the different curation process of the dataset.
Appendix G Post-training
Previous research report that LLMs that have undergone human-alignment procedures tend to have stronger political viewsSanturkar et al. (2023); Perez et al. (2023). Therefore, we also investigated the correlation between OLMO’s political leanings and the stance scores from the instruction-tuning dataset TULU , as well as the RLHF dataset UltraFeedback. However, no significant correlation was observed. This could be attributed to the small size of the documents, and only limited number of relevant documents retrieved from these datasets—only 15 out of the 32 cases had relevant documents in TULU, and just 10 cases had relevant documents in UltraFeedback. Prior research Feng et al. (2023) also suggests that the shift introduced by post-training is relatively small. We also explored the correlation between LLMs’ political leanings and that in their post-training data, but did not observe any significant correlation.
The key difference between OLMo-SFT (Supervised Fine-Tuning) and OLMo-Instruct lies in their fine-tuning objectives and intended uses. OLMo-SFT is fine-tuned for general language tasks using labeled data, using the TULU dataset Ivison et al. (2023). It is optimized for structured responses but isn’t specifically trained to follow user instructions. OLMo-instruct is further fine-tuned to follow human instructions, using the Ultrafeedback dataset Cui et al. (2023). It is optimized for handling detailed user instructions and conversational prompts, ideal for interactive and task-oriented use.
Appendix H Williams Test Results
Fig 12 includes a comprehensive overview of the Williams Test results of all LLMs.
Figure 12: Bootstrapped sample means and their 95% confidence intervals for each docket. Each bar represents the average stance score for a given case docket, while the error bars denote the 5th and 95th percentiles of the bootstrap distribution (based on repeatedly sampling 80% of the data).
Appendix I Example of a Retrieved Document
Fig 10 demonstrates the full text of a relevant document we retrieved from the pretraining dataset Dolma. The document is on case McDonald v. Chicago about the topic of gun control:
Example of a Retrieved Document In the run-up to the 2008 presidential election, many gun owners worried about the consequences of victory for Democrat candidate Barack Obama. Given Obama’s record as an Illinois state senator, where he stated his support for an all-out ban on handguns, among other gun control stances, pro-gun advocates were concerned that gun rights might suffer under an Obama presidential administration. After Obama’s election, gun sales reached a record pace as gun owners snatched up guns, particularly those that had been branded assault weapons under the defunct 1994 assault weapons ban, out of an apparent fear that Obama would crack down on gun ownership. The Obama presidency, however, had limited impact gun rights. When Obama was running for the Illinois state senate in 1996, the Independent Voters of Illinois, a Chicago-based non-profit, issued a questionnaire asking if candidates supported legislation to “ban the manufacture, sale, and possession of handguns,” to ”ban assault weapons” and to instate “mandatory waiting periods and background checks” for gun purchases. Obama answered yes on all three accounts. Obama also cosponsored legislation to limit handgun purchases to one per month. He also voted against letting people violate local weapons bans in cases of self-defense and stated his support for the District of Columbia’s handgun ban that was overturned by the U.S. Supreme Court in 2008. He also called it a “scandal” that President George W. Bush did not authorize a renewal of the Assault Weapons Ban. Just weeks after Obama’s inauguration in January 2009, attorney general Eric Holder announced at a press conference that the Obama administration would be seeking a renewal of the expired ban on assault weapons. “As President Obama indicated during the campaign, there are just a few gun-related changes that we would like to make, and among them would be to reinstitute the ban on the sale of assault weapons,” Holder said. U.S. Rep. Carolyn McCarthy, D-New York, introduced legislation to renew the ban. However, the legislation did not receive an endorsement from Obama. In the aftermath of a mass shooting in Tucson, Ariz., that wounded U.S. Rep. Gabrielle Giffords, Obama renewed his push for “common sense” measures to tighten gun regulations and close the so-called gun show loophole. While not specifically calling for new gun control measures, Obama recommended strengthening the National Instant Background Check system in place for gun purchases and rewarding states supplying the best data that would keep guns out of the hands of those the system is meant to weed out. Later, Obama directed the Department of Justice to begin talks about gun control, involving “all stakeholders” in the issue. The National Rifle Association declined an invitation to join the talks, with LaPierre saying there is little use in sitting down with people who have “dedicated their lives” to reducing gun rights. As the summer of 2011 ended, however, those talks had not led to recommendations by the Obama administration for new or tougher gun laws. One of the Obama administration’s few actions on the subject of guns has been to strengthen a 1975 law that requires gun dealers to report the sale of multiple handguns to the same buyer. The heightened regulation, which took effect in August 2011, requires gun dealers in the border states of California, Arizona, New Mexico and Texas to report the sale of multiple assault-style rifles, such as AK-47s and AR-15s. The story through much of his first term in office was a neutral one. Congress did not take up serious consideration of new gun control laws, nor did Obama ask them to. When Republicans regained control of the House of Representatives in the 2010 midterm, chances of far-reaching gun control laws being enacted were essentially squashed. Instead, Obama urged local, state, and federal authorities to stringently enforce existing gun control laws. In fact, the only two major gun-related laws enacted during the Obama administration’s first term actually expand the rights of gun owners. The first of these laws, which took effect in February 2012, allows people to openly carry legally owned guns in national parks. The law replaced a Ronald Reagan era policy that required guns to remain locked in glove compartments or trunks of private vehicles that enter national parks. The other law allows Amtrak passengers to carry guns in checked baggage; a reversal of a measure put in place by President George W. Bush in response to the terrorist attacks of Sept. 11, 2001.” Obama’s two nominations to the U.S. Supreme Court, Sonia Sotomayor, and Elena Kagan were considered likely to rule against gun owners on issues involving the Second Amendment. However, the appointees did not shift the balance of power on the court. The new justices replaced David H. Souter and John Paul Stevens, two justices who had consistently voted against an expansion of gun rights, including the monumental Heller decision in 2008 and McDonald decision in 2010. Earlier in his first term, Obama had expressed his express support for the Second Amendment. “If you’ve got a rifle, you’ve got a shotgun, you’ve got a gun in your house, I’m not taking it away. Alright?” he said. However, the legislation to overhaul gun control failed on April 17, 2013, when the Republican-controlled Senate rejected a measure banning assault-style weapons and expanding gun-buyer background checks. In January 2016, President Obama began his final year in office by going around the gridlocked Congress by issuing a set of executive orders intended to reduce gun violence. According to a White House Fact Sheet, the measures aimed to improve background checks on gun buyers, increase community safety, provide additional federal funding for mental health treatment, and advance the development of “smart gun” technology. During his eight years in office, President Barack Obama had to deal with more mass shootings than any of his predecessors, speaking to the nation on the subject of gun violence at least 14 times. In each address, Obama offered sympathy for the loved ones of the deceased victims and repeated his frustration with the Republican-controlled Congress to pass stronger gun control legislation. After each address, gun sales soared. In the end, however, Obama made little progress in advancing his “common-sense gun laws” at the federal government level — a fact he would later call one of the biggest regrets of his time as president. In 2015, Obama told the BBC that his inability to pass gun laws had been“the one area where I feel that I’ve been most frustrated and most stymied.

### v3

@@
 Abstract
-The increased adoption of Large Language Models (LLMs) and their potential to shape public opinion have sparked interest in assessing these models’ political leanings.
-Building on previous research that compared LLMs and human opinions and observed political bias in system responses, we take a step further to investigate the underlying causes of such biases by empirically examining how the values and biases embedded in training corpora shape model outputs.
-Specifically, we propose a method to quantitatively evaluate political leanings embedded in the large pretraining corpora.
+Recent works have shown that Large Language Models (LLMs) have a tendency to memorize patterns and biases present in their training data, raising important questions about how such memorized content influences model behavior.
+One such concern is the emergence of political bias in LLM outputs.
+In this paper, we investigate the extent to which LLMs’ political leanings reflect memorized patterns from their pretraining corpora.
+We propose a method to quantitatively evaluate political leanings embedded in the large pretraining corpora.
 Subsequently we investigate to whom are the LLMs’ political leanings more aligned with, their pretrainig corpora or the surveyed human opinions.
@@
 Our findings reveal that LLMs strongly reflect the political leanings in their training data, and no strong correlation is observed with their alignment to human opinions as expressed in surveys.
-These results underscore the importance of responsible curation of training data and the need for robust evaluation metrics to ensure LLMs’ alignment with human-centered values.
+These results underscore the importance of responsible curation of training data, and the methodology for auditing the memorization in LLMs to ensure human-AI alignment.
 […]
@@
 Unveiling Political Leanings of LLMs on U.S. Supreme Court Cases
-Shanshan Xu1, Santosh T.Y.S.S1,Yanai Elazar2,3 Quirin Vogel1, Barbara Plank4,5, Matthias Grabmair1 1Technical University of Munich, Germany 2Allen Institute for AI 3University of Washington 4IT University of Copenhagen, Denmark 5LMU Munich & Munich Center for Machine Learning (MCML)
+Shanshan Xu1, Santosh T.Y.S.S1,Yanai Elazar2,3 Quirin Vogel1, Barbara Plank4, Matthias Grabmair1 1Technical University of Munich, Germany 2Allen Institute for AI 3University of Washington 4LMU Munich & Munich Center for Machine Learning (MCML)
 1 Introduction
-Values and opinions expressed by large language models (LLMs) such as ChatGPT OpenAI (2023) have the potential to influence public perceptions and shape societal views.
-As such, it is essential to evaluate the political leanings of LLMs.
-Previous research (e.g., Perez et al.
-2023; Hartmann et al.
-2023) found a misalignment between LLM outputs and human opinions, and several identified a left-leaning bias of LLMs.
+LLMs derive their knowledge primarily from their pre-training data, which are typically composed of internet text.
+These sources, however, tend to overrepresent certain perspectives and ideologies, leading to biased training distributions Galeazzi et al.
+(2024).
+Previous work reveals that LLMs tend memorize parts of their training data Carlini et al.
+(2021).
+As a result, LLMs risk memorizing and reproducing these biases in downstream tasks, with potential societal consequences such as reinforcing political polarization or misrepresenting minority views Feng et al.
+(2023).
+While recent research has highlighted the presence of political bias in LLM outputs, the extent to which these biases stem from memorized content in pretraining data remains underexplored.
 Figure 1: Assessing the political leanings of LLMs, and comparing it with that in their training data, and of human respondents.
-LLMs derive their knowledge primarily from their pre-training data, which are typically composed of internet text.
-However, the structure of online content regularly skews toward overrepresenting specific viewpoints Galeazzi et al.
-(2024), often introducing biases into these models.
-Such biases can then propagate through model outputs, reinforcing existing societal biases Feng et al.
-(2023).
-Despite increasing awareness of these issues, to the best of our knowledge, no prior work has shown how to quantitatively analyze political bias in the content of large pre-training corpora and subsequently assess how well LLM outputs align with human opinions versus the biases in training data.
 To address this gap, we propose a pipeline to retrieve relevant documents from the pretraining corpora, then evaluate the political leanings expressed in these documents, and subsequently assess the alignment of political leanings in pretraining corpora with the responses generated by the LLM.
@@
 Hereafter, we refer to the dataset as SCOPE: Supreme COurt Case Political Evaluation.
-survey data on US Supreme Court cases from political studies, this paper examines the political leanings of eight LLMs and five open-source pre-training corpora, comparing them to human survey responses and Supreme Court rulings [^2] We will release the code and data upon publication..
-The main contributions of our work are threefold:
+survey data on US Supreme Court cases from political studies, this paper examines the political leanings of eight LLMs and five open-source pre-training corpora, comparing them to human survey responses and Supreme Court rulings.
+[^2] Our code and data is available at https://github.com/TUMLegalTech/scotus_alignment The main contributions of our work are threefold:
   • We conduct a quantitative analysis of political bias in large pre-training corpora by examining the political stance of the documents in the corpora.
@@
   • Our empirical findings indicate that LLMs exhibit significant alignment with their training corpora, yet we do not find strong alignment with human opinions.
-This highlights the critical need for responsible data curation and rigorous evaluation of LLM alignment.
+This highlights the critical need for methods to detect and mitigate memorized political content in LLMs.
+We advocate for more transparency in curating training data for LLMs.
 2 Background
-Evaluating LLM-Human Alignment
+2.1 LLMs and their pretraining corpra
+Existing studies have explored the impact of biases in training corpora on LLM behavior, primarily through second-stage controlled training setups such as continual pretraining Feng et al.
+(2023); Chalkidis and Brandl (2024).
+While continual pretraining can offer valuable insights into the causal links between training data and model outputs, these studies rarely applied to study LLMs’ behavior based on initial pretraining phase, where biases are fundamentally embedded.
+Additionally, it is also computationally expensive to conduct such extensive continual training experiments on initial phase.
+An alternative strategy involves investigating the correlation between biases in training corpora and those in model outputs Seshadri et al.
+(2024).
+Previously this approach has been underexploited, primarily due to the limited accessibility of large-scale pretraining datasets.
+Many commercial LLM providers (e.g., GPT-4 by OpenAI 2023 and Claude by Anthropic 2023) disclose minimal information about their training sets, not even corpus size or data source.
+With the growing call in the academic community for transparency and accessibility of LLM pretraining data Pistilli et al.
+(2023); McMillan-Major et al.
+(2024), several organizations have begun to make large-scale pretraining datasets publicly available, including RedPajama (Weber et al., 2024) and Dolma Soldaini et al.
+(2024).
+These initiatives are complemented by the development of APIs and analytical tooling platforms, such as WIMDB Elazar et al.
+(2024), which facilitate comprehensive analysis of the corpora.
+In this paper, we leverage WIMDB to analyze the political leanings in five publicly accessible corpora and subsequently evaluate how these leanings correlate with the outputs generated by various LLMs.
+2.2 Evaluating LLM-Human Alignment
 Recent research has increasingly focused on probing LLMs political opinions.
 Most approaches typically follow a two-stage process: (1) assessing an LLM’s political stance on specific topics, and (2) measuring how closely its responses align with human opinions.
-A common strategy for evaluating LLM opinions involves using political orientation tests (e.g., Political Compass Test333www.politicalcompass.org/test, as in Röttger et al.
+A common strategy for evaluating LLM opinions involves using political orientation tests (e.g., Political Compass Test,333www.politicalcompass.org/test as in Röttger et al.
 2024; Feng et al.
-2023) or survey questionnaires (e.g., PewResearch ATP444www.pewresearch.org/writing-survey-questions/, as in Santurkar et al.
+2023) or survey questionnaires (e.g., PewResearch ATP,444www.pewresearch.org/writing-survey-questions/ as in Santurkar et al.
 2023).
@@
 Moreover, political experts have annotated each case with thoughtfully chosen keywords, which facilitate our retrieval of relevant documents from large pretraining corpora, as detailed in Sec 4.1.3.
-LLMs and their pretraining corpra
-Existing studies have explored the impact of biases in training corpora on LLM behavior, primarily through second-stage controlled training setups such as continual pretraining Feng et al.
-(2023); Chalkidis and Brandl (2024).
-While continual pretraining can offer valuable insights into the causal links between training data and model outputs, these studies rarely applied to study LLMs’ behavior based on initial pretraining phase, where biases are fundamentally embedded.
-Additionally, it is also computationally expensive to conduct such extensive pre-training experiments on initial phase.
-An alternative strategy involves investigating the correlation between biases in training corpora and those in model outputs Seshadri et al.
-(2024).
-Previously this approach has been under exploited, primarily due to the limited accessibility of large-scale pretraining datasets.
-Many commercial LLM providers (e.g., GPT-4 by OpenAI 2023 and Claude by Anthropic 2023) disclose minimal information about their training sets, not even corpus size or data source.
-With the growing call in the academic community for transparency and accessibility of LLM pretraining data Pistilli et al.
-(2023); McMillan-Major et al.
-(2024), several organizations have begun to make large-scale pretraining datasets publicly available, including RedPajama (Weber et al., 2024) and Dolma Soldaini et al.
-(2024).
-These initiatives are complemented by the development of APIs and analytical tooling platforms, such as WIMDB Elazar et al.
-(2024), which facilitate comprehensive analysis of the corpora.
-In this paper, we leverage WIMDB to analyze the political leanings in five publicly accessible corpora and subsequently evaluate how these leanings correlate with the outputs generated by various LLMs.
 3 Experimental Setup
@@
 public.
-In this study, we use SCOPE Jessee et al.
-(2022) to prompt various LLMs to assess their political leanings and subsequently compare their alignment with surveyed human opinions and political leanings in their training data.
-The SCOPE dataset comprises of 32 cases, each represented by a binary-choice question asking respondents to express their views on the Court’s ruling as either supportive (pro) or opposing (opp).
+In this study, we use SCOPE to prompt various LLMs to assess their political leanings and subsequently compare their alignment with surveyed human opinions and political leanings in their training data.
+The SCOPE dataset comprises 32 cases, each represented by a binary-choice question asking respondents to express their views on the Court’s ruling as either supportive (pro) or opposing (opp).
 Fig 2 provides an example of a survey question.
@@
 BLOOMZ BLOOMZ-7b1 7B OSCAR*, The Pile*
-Table 1: Details of evaluated LLMs, along with the dataset that the model trained on.
+Table 1: Overview of evaluated LLMs, along with their pretraining dataset.
 * signifies that the model was not trained exactly on this dataset, due to filtering, using additional data, or the original data being private.
@@
 Tab 1 lists the corresponding pretraining corpora (when available) of the LLMs we investigated in this work.
-It is important to note that among the various LLM-corpora pairs we consider, only the OLMo-SFT and OLMo-Instruct models were trained directly on the pretraining corpus Dolma Soldaini et al.
+It is important to note that among the various pairs of LLM and their pretraining corpora we consider, only the OLMo-SFT and OLMo-Instruct models were trained directly on the pretraining corpus Dolma Soldaini et al.
 (2024).
@@
 (2024).
-Despite these discrepancies, we treat the documented corpora as reasonable proxies for analysis, as they represent the closest publicly available approximations of the actual training data for these models [^5] All models examined in this paper have undergone post-training, such as Supervised or Instruction Fine-Tuning, which may also influence the opinions in models outputs.
+Despite these discrepancies, we treat the documented corpora as reasonable proxies for analysis, as they represent the closest publicly available approximations of the actual training data for these models.
+[^5] All models examined in this paper have undergone post-training, such as Supervised or Instruction Fine-Tuning, which may also influence the opinions in models outputs.
 However, prior research Feng et al.
@@
 We also explored the correlation between LLMs’ political leanings and that in their post-training data, but did not observe any significant correlation.
-Further discussions can be found in App G..
+Further discussions can be found in App G.
 4 Methodology
-We employ a three-stage process to examine LLMs’ political leanings and further compare their alignment with surveyed human opinions and their pretraining corpora, when available.
-First, we introduce how we assess the political leanings of different entities [^6] We use entity to refer to either a group of surveyed respondents, Supreme Court justices, LLM-generated responses, or content within the training data..
+Figure 3: Extracting the Preference Distributions of the Pretraining Corpora.
+We employ a three-stage process to examine LLMs’ political leanings and compare their alignment with surveyed human opinions and their pretraining corpora, whenever available.
+First, we introduce how we assess the political leanings of different entities.
+[^6] We use entity to refer to either a group of surveyed respondents, Supreme Court justices, LLM-generated responses, or content within the training data.
 Next, we measure the political leanings alignment among them in Sec 4.2.
@@
 In this section, we outline the methodology used to extract the preference distribution of various entities.
-We divide the entities to three categories - the humans $D_{H}$, the LLMs $D_{M}$, and the pretraining corpora $D_{C}$.
-4.1.1 The Humans
-Under the category of the humans, we consider the preference distribution of four entities: $D_{H}=\{D_{\mathrm{pub}},D_{\mathrm{dem}},D_{\mathrm{rep}},D_{\mathrm{court}}\}$.
+We divide the entities to three categories - humans $D_{H}$, LLMs $D_{M}$, and pretraining corpora $D_{C}$.
+Figure 4: For each survey case in SCOPE, we created six different prompt templates, and we then sample five responses from each of the six prompt variations, yielding in a total of 30 responses per case per model.
+4.1.1 Humans
+Under the category of humans, we consider the preference distribution of four entities: $D_{H}=\{D_{\mathrm{pub}},D_{\mathrm{dem}},D_{\mathrm{rep}},D_{\mathrm{court}}\}$.
 Here, $D_{\mathrm{pub}}$ represents the preference distribution of the overall surveyed respondents, while $D_{\mathrm{dem}}$ and $D_{\mathrm{rep}}$ correspond to surveyed self-identified Democrat and Republican respondents, respectively; $D_{\mathrm{court}}$ represents the preference distribution of the Court.
@@
 For the $D_{\mathrm{court}}$, we fetch the judges’ votes from the Supreme Court Database Spaeth et al.
-(2024) [^7] http://scdb.wustl.edu/ , and then calculate $D_{\mathrm{court}}$ as the ratio of justices who agree (pro) / dissent (opo) with the majority decision.
-For $D_{pub},D_{\mathrm{dem}}$ and $D_{\mathrm{rep}}$ we calculate them as the ratios of {pro} versus {opo} to the court’s decisions among the respondents based on data retrieved from SCOPE.
-Figure 3: Extracting the Preference Distributions of the Pretraining Corpora.
-4.1.2 The LLMs
+(2024), [^7] http://scdb.wustl.edu/ and then calculate $D_{\mathrm{court}}$ as the ratio of justices who agree (pro) / dissent (opp) with the majority decision.
+For $D_{pub},D_{\mathrm{dem}}$ and $D_{\mathrm{rep}}$ we calculate them as the ratios of {pro} versus {opp} to the court’s decisions among the respondents based on data retrieved from SCOPE.
+4.1.2 LLMs
 Under the category of the LLMs $D_{M}$, we probe the political preferences of eight LLMs as listed in Tab 1.
 Following Scherrer et al.
-2023, for each survey case in SCOPE, we created six different prompt templates, and we then sample five responses from each of the six prompt variations from the LLMs at a temperature setting of 1, yielding in a total of 30 responses per case per model.
-Example prompt templates and detailed prompt creation process can be found in Fig 6 in App B.
-To map the LLM-generated answers to one of the given choice options, we employed an iterative, rule-based matching pipeline, as explained in App B.
-The preference distribution, denoted as $D_{m}=p_{m}(a_{j}\mid q_{i})$, reflects the ratio of support versus opposition to the court’s decision across the 30 generated responses for model $m$ on case $q_{i}$, illustrated in Fig 6 in App B.
-4.1.3 The Training Data
+2023, for each survey case in SCOPE, we created six different prompt templates, as illustrated in Fig 4.
+We then sample five responses from each of the six prompt variations from the LLMs at a temperature setting of 1, yielding in a total of 30 responses per case per model.
+The complete prompt templates and detailed prompt creation process can be found in 7(c) in App B.
+To map the LLM-generated answers to one of the given choice options, we employed an iterative, rule-based matching pipeline, as explained in App B, as illustrated in Fig 4.
+The preference distribution, denoted as $D_{m}=p_{m}(a_{j}\mid q_{i})$, reflects the ratio of support versus opposition to the court’s decision across the 30 generated responses for model $m$ on case $q_{i}$.
+4.1.3 Pretraining Data
 […]
@@
 For each of the 32 cases $q_{i}$ in the SCOPE survey, we compile a set of keywords $K_{i}$ to retrieve relevant documents $T_{i}$ from the pretraining corpora using the WIMDB API Elazar et al.
-(2024)- a tool designed to facilitate analysis of large-scale pretraining corpora.
+(2024), a tool designed to facilitate analysis of large-scale pretraining corpora.
 For example, in the case Baze v.
-Rees [^8] Baze v.
+Rees, [^8] Baze v.
 Rees, 553 U.S.
-35 (2008), addresses whether lethal injection for executions was constitutional or not., we use keywords such as [lethal injection; capital punishment; Baze; Rees], retrieving 206 documents from the Dolma corpus.
+35 (2008), addresses whether lethal injection for executions was constitutional or not.
+we use keywords such as [lethal injection; capital punishment; Baze; Rees] retrieving 206 documents from the Dolma corpus.
 Further details on keyword selection and retrieval statistics for each case are provided in App C.
@@
 (ii) Stance Score Evaluation
-We use Llama3-70B to assess the political stance $s_{i}^{j}$ of each retrieved document $t_{i}^{j}\in T_{i}$.
+We use zero-shot Llama3-70B to assess the political stance $s_{i}^{j}$ of each retrieved document $t_{i}^{j}\in T_{i}$.
 The model is prompted to evaluate the document’s level of support for the court’s decision on a Likert scale from 1 (strongly against) to 5 (strongly supportive).
-If a document is unrelated to the case’s political issue, the model is instructed to return ‘Not related.’
+If a document is unrelated to the case’s political issue, the model is instructed to return ‘Not related’.
+The complete prompt template we sed to evaluate the stance scores of the retrieved documents is available in Fig 8 in App B.
 (iii) Preference Distribution Estimation
@@
 To facilitate probabilistic interpretation, we transform $S_{i}$ from its original Likert scale to a probability measure $P_{i}$, which represents the likelihood that the document supports the court’s decision.
-Quality Assessment of Stance Detection To evaluate the reliability of Llama370B’s stance detection, we manually annotated a randomly selected sample of 80 retrieved documents.
+Quality Assessment of Stance Detection To evaluate the reliability of Llama3-70B’s stance detection, we manually annotated a randomly selected sample of 80 retrieved documents.
 We measure the agreement between human and model labels using Spearman’s rank correlation Spearman (1904).
@@
 The $p$-value associated to the Pearson coefficient quantifies statistical significance Kowalski (1972).
-4.3 Testing for Significance of Stronger Alignments
+4.3 Testing for Significance of Alignments
 Given an LLM $D_{m}$ and two human groups $D_{\mathrm{dem}}$ and $D_{\mathrm{rep}}$, we compute the alignments $\rho(D_{m},D_{\mathrm{dem}})$ and $\rho(D_{m},D_{\mathrm{rep}})$.
@@
 $K=1-\rho_{12}^{2}-\rho_{13}^{2}-\rho_{23}^{2}+2\rho_{12}\rho_{13}\rho_{23}\,.$
-Figure 4: Pearson Alignment. Cell $(i,j)$ represents the Pearson correlation $\rho$ of LLM $i$ to entity $j$. * shows p-value < 0.05, **∗ ∗ shows p-value < 0.001.
+Figure 5: Pearson Alignment.
+Cell $(i,j)$ represents the Pearson correlation $\rho$ of LLM $i$ to entity $j$.
+* shows p-value < 0.05, **∗ ∗ shows p-value < 0.001.
+random_1 stands for randomized values used as a baseline.
 Figure 6: The result of Williams significance tests, in each subfigure, where a colored cell in row $i$ (named on y-axis), col $j$ (named on x-axis) indicates that the LLM $m$ correlates significantly higher with entity $i$ than entity $j$, at a significance level of $0.05$.
@@
 Furthermore, all LLMs, with the exception of Bloomz and T0, display a significant positive correlation with the three training corpora: Dolma, RedPajama, and Oscar.
-This alignment may stem from the similar political leanings in these corpora and the models trained on them [^9] Fig 11 in App F presents the alignments between different training corpora and surveyed human opinions.
-.
+This alignment may stem from the similar political leanings in these corpora and the models trained on them.
+[^9] Fig 12 in App F presents the alignments between different training corpora and surveyed human opinions.
 In contrast, our findings indicate that there are generally no significant alignments between the LLMs’ outputs and surveyed human opinions.
@@
 At first glance, this might suggest that T0 is most aligned with human opinions among all entities.
-However, the significance test results in Fig 5(e) reveal inconsistencies.
+However, the significance test results in the Subfig (e) in Fig 6 reveal inconsistencies.
 While correlation between $(T0,E_{\text{pub}})$ is significantly stronger than the correlation between $(T0,E_{\text{court}})$, no such significant differences are found with other entities, such as with $E_{\text{dem}}$ or any of the training corpora.
@@
 Our research reveals that most LLMs exhibit alignment with their training corpora, yet not necessarily with the surveyed human opinions.
-Nonetheless, in the public discourse framework, attributing human characteristics to AI - known as anthropomorphizing — seems to be quite natural.
+Nonetheless, in the public discourse framework, attributing human characteristics to AI, also known as anthropomorphizing, seems to be quite natural.
 This tendency may lead to an over-reliance on AI, as users might confuse AI-generated responses for human beings, leading to excessive trust PAIR (2019).
@@
 By examining LLMs’ stances on political issues derived from U.S.
-Supreme Court cases, our results reveal a significant alignment between the models and their training corpora, yet no similarly strong alignment with human opinions as survyed is found.
-We call on the AI community to explore methods for rigorous evaluation of LLM alignment, and advocate for more transparent and collaborative strategies in curating training data for LLM.
+Supreme Court cases, our results reveal a significant alignment between the models and their training corpora, yet no similarly strong alignment with human opinions is found.
+These findings suggest that political bias in LLMs may be at least partly a result of memorization of biased content from pretraining corpora.
+We call on the AI community to explore methods for detecting, and mitigating memorized political bias in LLMs, and advocate for more transparent and collaborative strategies in curating training data for LLMs.
 Limitations
@@
 Acknowledgments
-We thank Rashid Haddad, Lukas Gosch, Valentin Hoffman, Max Prior and Yang Li for the helpful discussions.
+We thank Rashid Haddad, Lukas Gosch, Valentin Hoffman, Max Prior and Yang Li for the helpful discussions. BP is supported by the ERC Consolidator Grant DIALECT 101043235.
 Information about use of AI assistants
@@
 We collected all GPT responses in November 2024.
-Figure 6: LLM prompt
 Appendix B LLM Response Collection
@@
 As shown in Fig 9 , all sample means lie within their respective CIs, confirming the reliability of our estimates and quantifying their variability.
-docket Survey wave # doc fetched avg. length doc fetched avg. stance score # doc matched avg. length doc matched
-19-123 2021 59 1,346 3.40 113 11,111
-19-1257A 2021 124 1,369 2.27 165 5,414
-19-1257B 2021 384 1,148 4.42 441 4,900
-19-251 2021 338 909 4.13 396 2,784
-20-255 2021 520 978 4.23 577 1,739
-18-1259 2021 150 2,256 4.01 2,139 49,361
-19-783 2021 257 1,573 4.27 1,205 24,803
-20-512 2021 602 1,292 1.46 683 3,830
-20A87 2021 96 1,848 3.57 206 12,722
-20-107 2021 282 993 1.47 300 1,592
-19-422 2021 277 1,862 3.19 825 22,763
-20-18 2021 57 1,323 4.76 488 122,176
-17-1618 2020 87 1,513 4.75 102 4,025
-18-107 2020 405 1,246 4.87 562 5,803
-18-587 2020 2,005 1,047 4.31 2,216 2,311
-18-1195 2020 224 1,318 4.02 292 4,835
-19-431 2020 699 1,204 3.99 908 6,311
-18-1323 2020 185 1,712 3.60 253 7,368
-19-635 2020 878 1,181 1.56 1,214 6,384
-19-715 2020 456 1,216 1.51 693 12,316
-19-7 2020 1,047 1,159 4.19 1,205 2,920
-19-465 2020 814 1,040 3.63 895 2,015
-08-1224 2010 316 1,389 2.21 565 11,954
-08-1521 2010 1,570 2,380 4.47 5,903 32,346
-08-472 2010 108 1,165 2.62 172 13,837
-07-1428 2010 72 1,884 3.12 141 10,650
-07-21 2010 727 1,366 3.08 945 5,139
-08-205 2010 187 2,613 2.89 731 20,631
-07-5439 2010 927 1,458 2.25 1,360 5,799
-05-908 2010 471 1,802 2.73 974 15,752
-05-380 2010 402 2,012 3.53 983 12,542
-05-184 2010 56 2,419 3.17 176 20,370
-Table 3: Descriptive statistics of the documents retrieved from the Dolma dataset.
 Figure 10: Distribution of length of all the matched documents.
-Figure 11: Pearson Alignment. Cell $(i,j)$ represents the Pearson correlation $\rho$ of LLM $i$ to entity $j$. * shows *p* value < 0.05, **∗ ∗ shows p-value < 0.001.
 Appendix F Corpora-Human Alignment
@@
 The key difference between OLMo-SFT (Supervised Fine-Tuning) and OLMo-Instruct lies in their fine-tuning objectives and intended uses.
-OLMo-SFT is fine-tuned for general language tasks using labeled data, using the TULU dataset Ivison et al.
+OLMo-SFT is fine-tuned for general language tasks using labeled data, using the TULU datasetIvison* et al.
 (2023).
@@
 Fig 11 includes a comprehensive overview of the Williams Test results of all LLMs.
-Figure 12: Bootstrapped sample means and their 95% confidence intervals for each docket.
-Each bar represents the average stance score for a given case docket, while the error bars denote the 5th and 95th percentiles of the bootstrap distribution (based on repeatedly sampling 80% of the data).
 Appendix I Example of a Retrieved Document
@@
 In 2015, Obama told the BBC that his inability to pass gun laws had been“the one area where I feel that I’ve been most frustrated and most stymied.
+Figure 11: Bootstrapped sample means and their 95% confidence intervals for each docket.
+Each bar represents the average stance score for a given case docket, while the error bars denote the 5th and 95th percentiles of the bootstrap distribution (based on repeatedly sampling 80% of the data).
+Figure 12: Pearson Alignment. Cell $(i,j)$ represents the Pearson correlation $\rho$ of LLM $i$ to entity $j$. * shows *p* value < 0.05, **∗ ∗ shows p-value < 0.001.
+docket Survey wave # doc fetched avg. length doc fetched avg. stance score # doc matched avg. length doc matched
+19-123 2021 59 1,346 3.40 113 11,111
+19-1257A 2021 124 1,369 2.27 165 5,414
+19-1257B 2021 384 1,148 4.42 441 4,900
+19-251 2021 338 909 4.13 396 2,784
+20-255 2021 520 978 4.23 577 1,739
+18-1259 2021 150 2,256 4.01 2,139 49,361
+19-783 2021 257 1,573 4.27 1,205 24,803
+20-512 2021 602 1,292 1.46 683 3,830
+20A87 2021 96 1,848 3.57 206 12,722
+20-107 2021 282 993 1.47 300 1,592
+19-422 2021 277 1,862 3.19 825 22,763
+20-18 2021 57 1,323 4.76 488 122,176
+17-1618 2020 87 1,513 4.75 102 4,025
+18-107 2020 405 1,246 4.87 562 5,803
+18-587 2020 2,005 1,047 4.31 2,216 2,311
+18-1195 2020 224 1,318 4.02 292 4,835
+19-431 2020 699 1,204 3.99 908 6,311
+18-1323 2020 185 1,712 3.60 253 7,368
+19-635 2020 878 1,181 1.56 1,214 6,384
+19-715 2020 456 1,216 1.51 693 12,316
+19-7 2020 1,047 1,159 4.19 1,205 2,920
+19-465 2020 814 1,040 3.63 895 2,015
+08-1224 2010 316 1,389 2.21 565 11,954
+08-1521 2010 1,570 2,380 4.47 5,903 32,346
+08-472 2010 108 1,165 2.62 172 13,837
+07-1428 2010 72 1,884 3.12 141 10,650
+07-21 2010 727 1,366 3.08 945 5,139
+08-205 2010 187 2,613 2.89 731 20,631
+07-5439 2010 927 1,458 2.25 1,360 5,799
+05-908 2010 471 1,802 2.73 974 15,752
+05-380 2010 402 2,012 3.53 983 12,542
+05-184 2010 56 2,419 3.17 176 20,370
+Table 3: Descriptive statistics of the documents retrieved from the Dolma dataset.